OpenStudy (anonymous):

what value of x solves the following proportion? 9/6= x/8

OpenStudy (anonymous):

\[\large \frac{9}{6}=\frac{3}{2}\] and \[\large\frac{3}{2}=\frac{x}{8}\iff x=\frac{8\times 3}{2}\]

OpenStudy (anonymous):

6x=9*8 so x=(8*9)/6 x=12